Suppose you have a spherical balloon filled with air at room temperature and 1.0 atm pressure; its radius is 17cm . you take the balloon in an airplane, where the pressure is 0.78 atm. If the temperature is unchanged, what's the balloon's new radius?
Answer #1

At constant temperature,

P1/P2=(V2/V1)

or, P1/P2=(r2/r1)3

(1.0/0.78)=(r2